unknown66:

Given quadrilateral DEFG is an isosceles trapezoid. Find DE. DE=4x-2 DG=2x+5 GF=5x-3 FE=3x+2

jhonyy9:

 isosceles trapezoid mean lateral sides are equal

jhonyy9:

DE = FG

jhonyy9:

substitute the given expressions inside this equality and calcule the x value

jhonyy9:

DE = ?

jhonyy9:

please collaborate i m here to help you - courage please

unknown66:

so...4x-2=5x-3

unknown66:

right?

jhonyy9:

yes continue

jhonyy9:

how you isolate the x in a part of equation ?

jhonyy9:

are you there ????

unknown66:

yes..you then take -4

unknown66:

and (-) from 4x and 5x

jhonyy9:

4x-2=5x-3 how you isolate the x terms in a part of equation ?

jhonyy9:

subtract from both sides 4x so what will get ?

unknown66:

4x-4=0 and 5x-4=1

jhonyy9:

i ve said subtract 4X from both sides of equation

unknown66:

then you have -2=x-3

jhonyy9:

4x-2=5x-3 yes - and now how you get on the right side just x ?

jhonyy9:

-2 = x -3 isolate on the right side the x

unknown66:

you take -3 and (-) it on its self and -2...?

jhonyy9:

not exactly add to both sides 3 so what will get ?

jhonyy9:

-2 +3 = x -3 +3

jhonyy9:

x = ?

unknown66:

its 1=x

jhonyy9:

perfect and now how you get the DE ?

jhonyy9:

what you know about DE ?

jhonyy9:

DE = ?

unknown66:

DE=FG

jhonyy9:

yes but like expression - given in the text there up

unknown66:

DE=4X-2..?

jhonyy9:

yes sure and for x what value you ve get above ?

unknown66:

do you mean 1=x?

jhonyy9:

yes sure and now you know that x = 1 and DE = 4x - 2 how you get the DE = ?

unknown66:

I'm not sure

jhonyy9:

write it please

jhonyy9:

how you get the DE when x = 1 ?

jhonyy9:

???

jhonyy9:

this is the last step and very very easy

unknown66:

do you replace x with 1?

jhonyy9:

exactly

jhonyy9:

why you dont make so then you know it ?

unknown66:

4x-2 becomes 1-2....?

jhonyy9:

DE = 4x -2 x = 1 DE = ?

unknown66:

1

jhonyy9:

no substitute the x value in place of x inside this expression and calcule the DE

unknown66:

4x-2 x=1 puting it in 4x 1-2=1

jhonyy9:

DE = 4x -2 4*1 why equal for you 1 ?

jhonyy9:

1 substitute in place of x not in place of 4x

unknown66:

ohhh

jhonyy9:

ok ? and so will get DE = 4*1 -2 = ?

unknown66:

4 *1-2=2...?

jhonyy9:

yes - exactly and now you know that DE = 2 and given again that DE = FG how you can prove this ?

jhonyy9:

DE = 2 => FG = 2

jhonyy9:

FG = what expression above on the top ?

unknown66:

FG=5x-3

jhonyy9:

yes sure and you know that x = 1 so you need calcule the FG = ?

unknown66:

so 5*1--3

jhonyy9:

yes

jhonyy9:

FG = ?

jhonyy9:

5*1 -3 = ?

unknown66:

2

jhonyy9:

so we have DE = 2 and FG = 2 do you understand these now ?

unknown66:

yes

jhonyy9:

ok was my pleasure bye bye

unknown66:

bye thx